OpenStudy (unknown):

Which answer is equal to the quotient in the expression below? (2x3 - x2 - 3x) (x2 + x) A.2x - 3 B.2x2 - 3x C.2x2 + x D.2x - 3x

myininaya (myininaya):

i dont see a quotient

OpenStudy (unknown):

it's divided by

myininaya (myininaya):

are you doing (2x^3-x^2-3x)/(x^2+x)

OpenStudy (unknown):

never mind that one.. how about this Which answer is equal to the difference in the expression below? (2x3 - 6x2 + 2x - 2) - (-2x3 - 3x2 + 3x + 3) A.4x3 - 3x2 + x - 5 B.4x3 - 3x2 - 5 C.4x3 + 9x2 + x - 5 D.4x3 - 3x2 - x - 5

myininaya (myininaya):

\[\frac{2x^3-x^2-3x}{x^2+x}\]

myininaya (myininaya):

is this what you meant?

myininaya (myininaya):

you don't want to do this one?

myininaya (myininaya):

ok for this next one distribute you the negative in front of the parenthesis and then put like terms together

OpenStudy (unknown):

no i justed guessed on it, this was my next one. but i forgot how to set up subtraction

myininaya (myininaya):

if you have (2y-3x)-(-5y+7x) =2y-3x-(-5y)+(-7x) =2y-3x+5y-7x =2y+5y-3x-7x =7y-10x see i distribute the negative 1 to each term in the second paranthese then i grouped my like terms together let me see you try and i can correct where you went wrong

OpenStudy (unknown):

where are the ys coming from?

OpenStudy (unknown):

i have (2x^3-6x^2+2x-2)+(2x^3-3x^2+3x+3)

myininaya (myininaya):

that was an example

OpenStudy (unknown):

oh haa

myininaya (myininaya):

see that i got rid the grouping symbols after distributing the negative 1 you have (2x3 - 6x2 + 2x - 2) - (-2x3 - 3x2 + 3x + 3) =2x^3-6x^2+2x-2+2x^3+3x^2-3x-3

myininaya (myininaya):

not put your like terms togther

myininaya (myininaya):

now*

OpenStudy (unknown):

okay, now next step?

OpenStudy (unknown):

ohh alright.. so that's... my two 2x^3

OpenStudy (unknown):

do i add them and get 4x^3?

myininaya (myininaya):

ok! what about your x^2's?

OpenStudy (unknown):

well -6x^2+3x^2

myininaya (myininaya):

right

myininaya (myininaya):

-6+3=-3 so -6x^2+3x^2=

OpenStudy (unknown):

-3x^2?

myininaya (myininaya):

:) ok x's?

OpenStudy (unknown):

-x?

myininaya (myininaya):

:) and finally the constants?

OpenStudy (unknown):

+1?

myininaya (myininaya):

-2-3 is not 1 2+3=5 so -(2+3)=-(5)=-5 -(2+3)=-2+(-3)=-2-3=-5

OpenStudy (unknown):

-5?

myininaya (myininaya):

yes so D is the answer

OpenStudy (unknown):

woop woop. haha thank youuu. :)

myininaya (myininaya):

np